### v1

# Pseudo-Autoregressive Neural Codec Language Models for Efficient Zero-Shot Text-to-Speech Synthesis

Abstract.
Recent zero-shot text-to-speech (TTS) systems face a common dilemma: autoregressive (AR) models suffer from slow generation and lack duration controllability, while non-autoregressive (NAR) models lack temporal modeling and typically require complex designs. In this paper, we introduce a novel pseudo-autoregressive (PAR) codec language modeling approach that unifies AR and NAR modeling. Combining explicit temporal modeling from AR with parallel generation from NAR, PAR generates dynamic-length spans at fixed time steps. Building on PAR, we propose PALLE, a two-stage TTS system that leverages PAR for initial generation followed by NAR refinement. In the first stage, PAR progressively generates speech tokens along the time dimension, with each step predicting all positions in parallel but only retaining the left-most span. In the second stage, low-confidence tokens are iteratively refined in parallel, leveraging the global contextual information. Experiments demonstrate that PALLE, trained on LibriTTS, outperforms state-of-the-art systems trained on large-scale data, including F5-TTS, E2-TTS, and MaskGCT, on the LibriSpeech test-clean set in terms of speech quality, speaker similarity, and intelligibility, while achieving up to ten times faster inference speed. Audio samples are available at https://anonymous-palle.github.io.
Zero-Shot Text-to-Speech, Pseudo-Autoregressive, Masked Language Modeling, Neural Codec Language Model
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: arXiv Preprint; April, 2025; arXiv.orgccs: Computing methodologiesccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Computing methodologies Natural language generation
Figure 1. Generation procedures of three paradigms, where generation proceeds top to bottom, and each row represents one step. (Left) AR generates tokens step by step from left to right until the [EOS] token. Grouped AR can predict multiple tokens per step. (Middle) NAR predicts all masked positions simultaneously in a temporally unordered manner, allowing partial token retention while remasking others for regeneration. (Right) PAR predicts all masked positions in parallel but commits only the leftmost span at each step, enforcing generation to follow the temporal order of speech.
[…]
A diagram showing generation procedures across AR, NAR, and PAR paradigms.
1. Introduction
Text-to-speech (TTS) synthesis has made remarkable progress, driven by generative methods (Shen et al., 2018; Li et al., 2019; Kim et al., 2021; Ren et al., 2021; Jeong et al., 2021; Wang et al., 2023) and the expansion of available data (Guo et al., 2021; Ma et al., 2024; Chen et al., 2021; Yang et al., 2024c; Kang et al., 2024; He et al., 2024). These innovations have enabled TTS systems to achieve human-level parity in terms of naturalness and intelligibility, in both pre-defined specific speakers (Tan et al., 2024) and zero-shot scenarios (Chen et al., 2024a).
Modern zero-shot TTS models can be broadly divided into two paradigms: autoregressive (AR) (Wang et al., 2023; Chen et al., 2024a; Meng et al., 2024; Wang et al., 2025; Yang et al., 2024a; Du et al., 2024a, c; Peng et al., 2024; Lajszczak et al., 2024; Anastassiou et al., 2024; Kharitonov et al., 2023; Kim et al., 2024; Liu et al., 2024; Sheng et al., 2025; Guo et al., 2024) [^1] For hybrid systems like VALL-E, we regard it as AR, as primary tokens are generated autoregressively. and non-autoregressive (NAR) (Shen et al., 2024; Ju et al., 2024; Le et al., 2023; Eskimez et al., 2024; Liu et al., 2025; Gao et al., 2023; Chen et al., 2024b; Wang et al., 2024; Casanova et al., 2022; Jiang et al., 2023, 2024, 2025; Lee et al., 2025). AR models sequentially generate speech from left to right, conditioning each step on previous outputs. Despite strong temporal grounding, AR models face fundamental limitations in generation efficiency and contextual modeling. Specifically, the fixed prediction size per step results in slow inference, with generation steps tightly coupled to the frame rate of output units. Additionally, the strictly unidirectional modeling restricts access to left context only and lacks future or global information, leading to exposure bias and posing challenges to robustness (Song et al., 2024a; Han et al., 2024; Du et al., 2024b). Techniques like grouped modeling (Chen et al., 2024a) and speculative decoding (Li et al., 2024) improve efficiency, but generation steps still scale linearly with synthesized speech duration. In contrast, NAR models enable fast inference through parallel generation. NAR TTS methods leverage denoising diffusion (Ju et al., 2024), flow matching (Eskimez et al., 2024; Chen et al., 2024b), GAN (Jiang et al., 2024), and masked generative modeling (Wang et al., 2024), drawing inspiration from advances in image (Chang et al., 2022) and language (Devlin et al., 2019) generation. Recent advances (Eskimez et al., 2024; Wang et al., 2024) demonstrate that text-to-speech alignment can be achieved without phoneme-level duration prediction or explicit alignment supervision. E2-TTS (Eskimez et al., 2024) and F5-TTS (Chen et al., 2024b) simply pad the character sequence with filler tokens to match the linearly predicted or user-specified speech length. MaskGCT (Wang et al., 2024) concatenates the tokenized text and speech sequences, adopts mask-and-predict training, and performs iterative parallel decoding during inference. However, despite their efficiency and simplicity, these methods lack temporal modeling and generate speech in a temporally unordered manner, resulting in slow convergence (Eskimez et al., 2024; Chen et al., 2024b), alignment inaccuracies (Eskimez et al., 2024), and degraded intelligibility (Wang et al., 2024).
To combine the strengths of AR and NAR paradigms while mitigating their limitations, we propose a novel pseudo-autoregressive (PAR) codec language modeling paradigm that unifies AR and NAR modeling by combining temporal modeling from AR with parallel generation from NAR. PAR introduces a soft temporal inductive bias into an NAR structure, specifically a bidirectional masked generative transformer, by enforcing span-level causal ordering through progressive commitment. Local spans are generated simultaneously, with earlier ones committed before later ones, creating a span-level temporal flow that aligns with the structure of natural speech. Unlike grouped AR methods (Chen et al., 2024a), which generate fixed-length spans with dynamic time steps, PAR generates dynamic-length spans with fixed time steps. Building on PAR, we propose PALLE, a two-stage TTS system that performs PAR generation followed by confidence-guided refinement. In the first stage, PALLE progressively commits spans along the temporal dimension, generating a fixed portion of the target sequence at each step, resulting in constant inference steps regardless of target speech duration. In the second stage, low-confidence tokens are refined in parallel based on global context, requiring only a few steps to correct accumulated errors and yielding huge performance improvements.
Experiments demonstrate that PALLE, trained on LibriTTS, outperforms state-of-the-art systems trained on large-scale data, including F5-TTS, E2-TTS, and MaskGCT, on the LibriSpeech test-clean set in terms of speech quality, speaker similarity, and intelligibility, with inference speed up to 10$\times$ faster.
Our contributions can be summarized as:
  • We propose and formulate a novel PAR codec language modeling paradigm for zero-shot TTS, which unifies AR and NAR paradigms by combining temporal modeling with parallel generation, offering new insights into generation algorithm design.
  • We instantiate AR, NAR, and PAR modeling paradigms under controlled conditions, and provide both empirical and theoretical indications of the effectiveness of PAR language modeling.
  • We introduce a new zero-shot TTS framework, namely PALLE, which employs PAR for initial generation followed by NAR confidence-guided refinement.
  • We achieve a breakthrough in zero-shot TTS performance and efficiency, with the small-scale trained PALLE outperforming large-scale state-of-the-art systems while enabling up to 10$\times$ faster inference.
2. Related work
2.1. Zero-Shot TTS
Zero-shot TTS enables speech synthesis for unseen speakers by capturing the timbre, prosody, and style from short speech prompts. Benefiting from in-context learning (ICL), the performance of zero-shot TTS has improved significantly, reaching human-parity quality (Ju et al., 2024; Chen et al., 2024a). Recent approaches primarily fall into three categories: codec language modeling (Wang et al., 2023; Chen et al., 2024a; Wang et al., 2024), diffusion/flow-based methods (Shen et al., 2024; Ju et al., 2024; Le et al., 2023; Eskimez et al., 2024; Chen et al., 2024b), and hybrid approaches (Du et al., 2024a, c). Hybrid systems involve a text-to-codec language model with a codec-to-mel model based on diffusion or flow matching, disentangling complex speech components to improve zero-shot generalization.
Figure 2. Overview of PALLE, a two-stage codec language model for zero-shot TTS, comprising the following components: (1) a text tokenizer that converts text into subword tokens; (2) a speech tokenizer that encodes speech into a sequence of discrete speech tokens; (3) two masked generative transformer with shared architecture, where the first maps text tokens to speech tokens and the second refines generated speech tokens; and (4) a speech detokenizer with a built-in vocoder that converts the speech token sequence into waveform.
[…]
An overview diagram of PALLE.
2.2. Discrete Speech Representation
Discrete speech representations (Yang et al., 2024b) are typically categorized into acoustic tokens and semantic tokens. Acoustic tokens, produced by neural codec models (Défossez et al., 2023; Kumar et al., 2023; Zeghidour et al., 2022; Ji et al., 2025; Parker et al., 2024; Wu et al., 2024; Du et al., 2024d), aim to reconstruct speech with high fidelity, capturing fine-grained acoustic details. In contrast, semantic tokens are quantized from embeddings of self-supervised learning (SSL) models (Baevski et al., 2020a, b; Hsu et al., 2021; Chen et al., 2022), which are trained for discrimination or masked prediction, and preserve the linguistic content of speech. Recent methods have explored supervised semantic tokens (Du et al., 2024a, c), trained with automatic speech recognition (ASR) loss (Graves et al., 2006, 2013), making them more robust to noise and better suited for language modeling.
2.3. Neural Codec Language Modeling
Neural codec language modeling has emerged as a powerful paradigm for generative speech modeling. Built upon discrete speech representations, language modeling approaches are employed to predict speech tokens, which are then decoded into waveforms by a speech detokenizer. Two main paradigms have been explored: AR and NAR modeling. AR models (Wang et al., 2023; Chen et al., 2024a; Yang et al., 2024a; Du et al., 2024a, c; Kharitonov et al., 2023; Lajszczak et al., 2024; Sheng et al., 2025) generate tokens sequentially from left to right, providing strong temporal grounding but suffering from slow inference and limited context modeling. SyncSpeech (Sheng et al., 2025) is an AR masked generative model that incorporates temporal modeling via a causal attention mask and token-level duration prediction, aiming for streaming generation from partial text input. Both Grouped AR and SyncSpeech can predict multiple tokens at one time step, but the number of generation steps scales linearly with the input text, and only the previous context is considered during inference. In contrast, PAR generates the output sequence with fixed time steps while considering the global context. NAR masked generative models (Wang et al., 2024) improve generation efficiency by iteratively infilling masked positions using bidirectional context. MaskGCT (Wang et al., 2024) predicts the total speech length and performs temporally unordered prediction, relying solely on confidence-based scheduling, which may cause robustness issues in high-entropy regions such as fast transitions or expressive prosody.
3. Pseudo-Autoregressive Language Modeling
We propose a pseudo-autoregressive (PAR) codec language modeling approach for zero-shot TTS that unifies NAR and AR modeling. As illustrated in Fig. 1, PAR combines the parallel prediction from NAR with the temporal modeling from AR, enabling structured parallelism that follows the nature temporal order of speech.
Formulation Given a speech sample $\mathbf{y}$ and its corresponding tokenized text transcription $\mathbf{x}$, a pretrained speech tokenizer is used to encode the speech sample into discrete speech tokens $\mathbf{c}$ of downsampled length $T$, denoted as $\mathbf{c}=\mathrm{Encode}_{\text{spch}}(\mathbf{y})$.
The PAR model is built upon a bidirectional masked generative transformer (Chang et al., 2022). During training, the model is optimized to maximize the likelihood of the speech tokens $\mathbf{c}$ conditioned on the text content $\mathbf{x}$ using cross-entropy loss. At each training step, a contiguous span mask is applied to $\mathbf{c}$ from a random starting position to the end:
(1) $\mathbf{m}=[\underbrace{0,0,\ldots,0}_{T-\ell},\underbrace{1,1,\ldots,1}_{\ell}],$
where $\ell$ denotes the length of the masked span. The model is optimized to predict the leftmost span of the masked portion $\mathbf{m}\odot\mathbf{c}$ with length $k=\lfloor rT\rfloor$, where $r\in(0,1)$ is a fixed ratio. This prediction is conditioned on the unmasked portion $(1-\mathbf{m})\odot\mathbf{c}$ and the text content $\mathbf{x}$, corresponding to maximizing the following objective:
(2) $\arg\max_{\theta}\;p\left([\mathbf{m}\odot\mathbf{c}]_{0:k}\mid(1-\mathbf{m})\odot\mathbf{c},\mathbf{x};\theta\right),$
where $[\,\cdot\,]_{0:k}$ denotes the first $k$ elements, and $\theta$ represents the model parameters. The mask starting position is dynamically sampled during training, enabling the model to predict the next span along the time dimension based on varying amounts of preceding speech context. Note that all self-attention layers use full attention.
During inference, given a tokenized text $\mathbf{x}^{\text{gen}}$ to be synthesized, a speech prompt $\mathbf{y}^{\text{ref}}$ from an unseen speaker, and its corresponding tokenized text $\mathbf{x}^{\text{ref}}$, the speech tokens $\mathbf{c}^{\text{ref}}=[c^{\prime}_{0},c^{\prime}_{1},\ldots,c^{\prime}_{T^{\text{ref}}-1}]$ of length $T^{\text{ref}}$. The target speech token length $T^{\text{gen}}$ can be determined arbitrarily or estimated. Inference begins by initializing an extended speech token sequence:
(3) $\mathbf{c}^{\text{ext}}_{(0)}=[c^{\prime}_{0},c^{\prime}_{1},\ldots,c^{\prime}_{T^{\text{ref}}-1},\underbrace{0,0,\ldots,0}_{T^{\text{gen}}-T^{\text{ref}}}],$
where the appended zeros serve as placeholders for the speech tokens to be generated. At each decoding step $t$, the model predicts all tokens in parallel, but retains only the leftmost $k^{\prime}=\min(\lfloor r^{\prime}T^{\text{gen}}\rfloor,N_{\text{left}})$ tokens, where $r^{\prime}\in(0,1)$ is a fixed ratio and $N_{\text{left}}$ is the number of tokens yet to be generated. Predictions are conditioned on the text prompt $\mathbf{x}^{\text{ref}}$, the target text $\mathbf{x}^{\text{gen}}$, and the current extended speech token sequence $\mathbf{c}^{\text{ext}}_{(t)}$. A binary mask is constructed as follows:
(4) $\mathbf{m}_{(t)}=[\underbrace{0,0,\ldots,0}_{T^{\text{ref}}\;+\;t\times k^{\prime}},\underbrace{1,1,\ldots,1}_{k^{\prime}},0,0,\ldots,0],$
The extended speech token sequence is updated according to the following iterative rule:
$\mathbf{c}^{\text{ext}}_{(t+1)}$ $=(1-\mathbf{m}_{(t)})\odot\mathbf{c}^{\text{ext}}_{(t)}$
(5) $+\mathbf{m}_{(t)}\odot\arg\max_{\mathbf{c}^{\text{ext}}_{(t+1)}}p\left(\mathbf{c}^{\text{ext}}_{(t+1)}\mid\mathbf{c}^{\text{ext}}_{(t)},\mathbf{x}^{\text{ref}},\mathbf{x}^{\text{gen}};\theta\right),$
where the first part contains the tokens of the speech prompt and previously generated spans, and the second part contains the currently generated tokens. This generation and selectively updating continue until all the tokens are generated. The final speech tokens $\mathbf{c}^{\text{ext}}_{(\infty)}$ are converted into waveform $\hat{\mathbf{y}}$ using a pretrained speech detokenizer and vocoder, jointly denoted as $\hat{\mathbf{y}}=\mathrm{Decode}_{\text{spch}}(\mathbf{c}^{\text{ext}}_{(\infty)})$.
Discussion PAR is defined as a general language modeling paradigm that does not impose constraints on implementation details such as the fusion of text and speech tokens. It accommodates both temporal concatenation, as in MaskGCT (Wang et al., 2024), and feature-dimension fusion, where text and speech tokens are first padded to the same length, then independently embedded and concatenated along the feature dimension, as in E2 TTS (Eskimez et al., 2024).
Figure 3. Illustration of the two-stage PALLE framework. (a) The shared architecture: a bidirectional masked generative transformer. (b) Stage one: the model predicts all token positions in parallel but retains only the leftmost span at each step. (c) Stage two: the model refines the initial speech tokens, where low-confidence tokens are re-masked and regenerated using contextual information.
[…]
An illustration diagram of PALLE framework.
4. PALLE
4.1. Architecture
As illustrated in Fig. 2, given the text to be synthesized, a BPE-based text tokenizer converts the text into subword tokens, which are then processed by two masked generative transformers with a shared architecture. The first stage generates speech tokens with a PAR model, and the second stage refines the speech tokens with an NAR model. Finally, a speech detokenizer with a built-in vocoder converts the refined speech tokens into the waveform.
Masked Generative Language Models As depicted in Fig. 3(a), both stages in PALLE adopt a bidirectional masked generative transformer (Chang et al., 2022) as the backbone. The model processes two input sequences of equal length, a padded tokenized text sequence and a masked speech token sequence. The text embeddings are further processed by a ConvNeXt V2 block (Woo et al., 2023), which provides strong temporal modeling capabilities, making the text representation better aligned with the speech sequence. A sinusoidal positional encoding is added for both sequences with a learnable scaling factor and dropout. The two sequences are concatenated along the feature dimension and passed through a linear projection layer. A convolutional positional embedding (Le et al., 2023) is then added to the projected features to encode temporal structure. The resulting sequence is fed into a bidirectional transformer that attends to the full context.
Speech Tokenizer and Detokenizer We use the pretrained S3Tokenizer from CosyVoice 2 (Du et al., 2024a) to extract discrete semantic tokens from the input waveform at a rate of 25 Hz. This model is a fine-tuned version of the SenseVoice-Large (An et al., 2024) ASR model, trained on a large multilingual speech dataset for robust speech understanding. By leveraging ASR loss during training, the S3Tokenizer extracts semantic information, enabling implicit denoising and speaker disentanglement (Song et al., 2024b). We adopt the pretrained conditional flow matching (CFM) model from CosyVoice 2 (Du et al., 2024c) to reconstruct mel spectrograms from discrete speech tokens. The generated mel spectrogram is then converted into the waveform using the pretrained HiFi-GAN (Kong et al., 2020) vocoder from CosyVoice 2 (Du et al., 2024c).
4.2. Training: Conditional Codec Masked Language Modeling
For simplicity, we reuse the symbol $\mathbf{c}=[c_{0},c_{1},\ldots,c_{T-1}]$ to denote a sequence of speech tokens obtained from a speech sample, and $\mathbf{x}=[x_{0},x_{1},\ldots,x_{L-1}]$ of length $L$ to represent its tokenized transcript.
Stage One The first stage of PALLE implements PAR with the modality fusion method used in E2 TTS (Eskimez et al., 2024). As depicted in Fig. 3(b), the text token sequence $\mathbf{x}$ is padded with the filler token [PAD] to match the length of the speech token sequence $\mathbf{c}$, enabling fusion in the feature dimension after embedding. The padded text sequence $\mathbf{x}^{\text{ext}}$ is defined as:
(6) $\mathbf{x}^{\text{ext}}=[x_{0},x_{1},\ldots,x_{L-1},\underbrace{\texttt{[PAD]},\texttt{[PAD]},\ldots,\texttt{[PAD]}}_{T-L}].$
At each step, a contiguous span of speech tokens in $\mathbf{c}$ is randomly masked using the binary mask $\mathbf{m^{\prime}}$, which extends Eq. 1 by keeping the first 30% speech tokens unmasked as a prompt. The starting position $s$ is sampled from a uniform distribution:
(7) $s\sim\mathcal{U}\left\{\lfloor 0.3T\rfloor,\lfloor 0.3T\rfloor+1,\dots,T-\lfloor 0.1T\rfloor-1\right\},$
where the upper bound of the range ensures that at least the final 10% of the sequence could be used for training. Based on $s$, the binary mask $\mathbf{m^{\prime}}$ is defined as:
(8) $\mathbf{m^{\prime}}=[\underbrace{0,0,\ldots,0}_{s},\underbrace{1,1,\ldots,1}_{T-s}].$
The model is optimized to predict the leftmost span of size $k=\left\lfloor 0.1T\right\rfloor$ from the masked speech token sequence, denoted as $[\mathbf{m^{\prime}}\odot\mathbf{c}]_{0:k}$, conditioned on the unmasked portion $(1-\mathbf{m^{\prime}})\odot\mathbf{c}$ and the padded text sequence $\mathbf{x}^{\text{ext}}$. The training objective is:
(9) $\arg\max_{\theta}\;p\left([\mathbf{m^{\prime}}\odot\mathbf{c}]_{0:k}\mid(1-\mathbf{m^{\prime}})\odot\mathbf{c},\mathbf{x}^{\text{ext}};\theta\right).$
Stage Two Similar to stage one, $\mathbf{x}$ is padded with the filler token [PAD] to match the length of the speech token sequence $\mathbf{c}$. As illustrated in Fig. 3(c), speech tokens beyond $\lfloor 0.3T\rfloor$ are independently masked with probability $p=0.1$, forming a binary mask $\mathbf{m^{\prime\prime}}=\{0\}^{\lfloor 0.3T\rfloor}\oplus\{0,1\}^{T-\lfloor 0.3T\rfloor}$, where $\oplus$ means cancatenation and 1 indicating masked positions. The model is optimized to predict the masked speech tokens $\mathbf{m^{\prime\prime}}\odot\mathbf{c}$, conditioned on the unmasked speech tokens $(1-\mathbf{m^{\prime\prime}})\odot\mathbf{c}$ and the padded text sequence $\mathbf{x}^{\text{ext}}$, maximizing the following objective:
(10) $\arg\max_{\theta}\;p\left(\mathbf{m^{\prime\prime}}\odot\mathbf{c}\mid(1-\mathbf{m^{\prime\prime}})\odot\mathbf{c},\mathbf{x}^{\text{ext}};\theta\right).$
4.3. Inference: In-Context Learning Via Prompting
We reuse the symbol $\mathbf{x}^{\text{gen}}=[x^{g}_{0},x^{g}_{1},\ldots,x^{g}_{L^{\text{gen}}-1}]$ of length $L^{\text{gen}}$ to represent the tokenized text to be synthesized, $\mathbf{y}^{\text{ref}}$ as the speech prompt, $\mathbf{x}^{\text{ref}}=[x^{r}_{0},x^{r}_{1},\ldots,x^{r}_{L^{\text{ref}}-1}]$ of length $L^{\text{ref}}$ as the tokenized text prompt, and $\mathbf{c}^{\text{ref}}=[c^{r}_{0},c^{r}_{1},\ldots,c^{r}_{T^{\text{ref}}-1}]$ of length $T^{\text{ref}}$ as the corresponding speech tokens from $\mathbf{y}^{\text{ref}}$.
Stage One The target speech token length $T^{\text{gen}}$ is either predefined or estimated linearly as in (Chen et al., 2024b):
(11) $T^{\text{gen}}=T^{\text{ref}}\times(1+\frac{L^{\text{gen}}}{L^{\text{ref}}}).$
The prompt speech token sequence $\mathbf{c}^{\text{ref}}$ is padded with the masked token [MASK] to match $T^{\text{gen}}$:
(12) $\mathbf{c}^{\text{ext}}=[c^{r}_{0},c^{r}_{1},\ldots,c^{r}_{T^{\text{ref}}-1},\underbrace{\texttt{[MASK]},\texttt{[MASK]},\ldots,\texttt{[MASK]}}_{T^{\text{gen}}-T^{\text{ref}}}].$
Meanwhile, the tokenized text sequences $\mathbf{x}^{\text{ref}}$ and $\mathbf{x}^{\text{gen}}$ are concatenated and padded with the filler token [PAD] to reach $T^{\text{gen}}$:
(13) $\mathbf{x}^{\text{ext}}=[x^{r}_{0},\ldots,x^{r}_{L^{\text{ref}}-1},x^{g}_{0},\ldots,x^{g}_{L^{\text{gen}}-1},\underbrace{\texttt{[PAD]},\ldots,\texttt{[PAD]}}_{T^{\text{gen}}-(L^{\text{ref}}+L^{\text{gen}})}].$
The model progressively generates speech tokens along the time dimension. At each step $t$, the model predicts all tokens in parallel, retaining only the leftmost span of size $k^{\prime}=\min(\left\lfloor r^{\prime}T\right\rfloor,N_{\text{left}})$, where $r^{\prime}\in(0,1)$ is a fixed ratio, and $N_{\text{left}}$ is the number of tokens yet to be generated . Predictions are conditioned on the current extended speech and text token sequences, $\mathbf{c}^{\text{ext}}_{(t)}$ and $\mathbf{x}^{\text{ext}}$. A binary mask $\mathbf{m}_{(t)}$ is constructed as in Eq. 4, and the extended speech token sequence is updated iteratively, as in Eq.3:
$\mathbf{c}^{\text{ext}}_{(t+1)}$ $=(1-\mathbf{m}_{(t)})\odot\mathbf{c}^{\text{ext}}_{(t)}$
(14) $+\mathbf{m}_{(t)}\odot\arg\max_{\mathbf{c}^{\text{ext}}_{(t+1)}}p\left(\mathbf{c}^{\text{ext}}_{(t+1)}\mid\mathbf{c}^{\text{ext}}_{(t)},\mathbf{x}^{\text{ext}};\theta\right).$
Stage Two In this stage, the model iteratively refines the initial generation $\mathbf{c^{\prime}}^{\text{ext}}_{(0)}=\mathbf{c}^{\text{ext}}_{(\infty)}$ from stage one, by re-masking and re-predicting low-confidence tokens. At each step $t$, the model produces a probability matrix $\mathbf{P}_{(t)}$, where each entry $P_{(t),n}$ denotes the predicted probability of class $n$ out of $N$ speech token classes, conditioned on $\mathbf{c^{\prime}}^{\text{ext}}_{(t)}$ and $\mathbf{x}^{\text{ext}}$:
(15) $\mathbf{P}_{(t)}=p\left(\mathbf{c^{\prime}}^{\text{ext}}_{(t+1)}\mid\mathbf{c^{\prime}}^{\text{ext}}_{(t)},\mathbf{x}^{\text{ext}};\theta\right)\in\mathbb{R}^{T^{\text{gen}}\times N}.$
The confidence score matrix $\mathbf{C}_{(t)}\in\mathbb{R}^{T^{\text{gen}}}$ is defined as the negative min-entropy of the corresponding distribution:
(16) $\mathbf{C}_{(t)}=\log\mathbf{P}^{\text{max}}_{(t)}[:,n],$
where $\mathbf{P}^{\text{max}}_{(t)}$ is the maximum probabilities in the predicted distributions. We rank the confidence scores and select those tokens with the lowest confidence for re-masking using a predefined quantile $\gamma$. A binary mask $\mathbf{m}_{(t)}\in\{0,1\}^{T^{\text{gen}}}$ is constructed, where 1 marks the selected positions for masking. The extended speech token sequence $\mathbf{c^{\prime}}^{\text{ext}}_{(t)}$ is updated iteratively as in Eq. 4.3. To prevent repeated refinement at the same position, the confidence scores of updated tokens are set to 1 permanently. The final speech tokens $\mathbf{c^{\prime}}^{\text{ext}}_{(\infty)}$ are converted into waveform $\hat{\mathbf{y}}$.
Table 1. Objective performance comparison on continuation and cross-sentence zero-shot speech synthesis tasks. Bold highlights the best result, while underlined marks the second-best. Metrics not reported in the original papers are calculated using the open-source checkpoints. GT duration denotes the results using the ground-truth total speech duration. For Continuation task, NAR systems without GT duration are excluded, as linear duration estimation requires forced alignment to obtain the transcription of the first three seconds of speech.
System Speech Tokenizer Training Data Continuation Cross-Sentence
[…]
Ground Truth - - 2.14 2.15 0.905 2.14 2.15 0.779 -
Ground Truth (EnCodec) - - - 2.33 0.823 - 2.33 0.715 -
Ground Truth (S3Tokenizer v2 25Hz) - - 2.69 3.16 0.793 2.94 3.51 0.743 -
Trained on Large-Scale Dataset
VALL-E (Wang et al., 2023) EnCodec (Défossez et al., 2023) Librilight - 3.80 0.773 - 5.90 0.633 0.73
VALL-E  (Chen et al., 2025) EnCodec (Défossez et al., 2023) Libriheavy - 3.00 0.770 - 3.80 0.630 0.73
E2 TTS (32 NFE) - Emilia (EN + ZH) - - - 2.55 2.92 0.756 0.68
F5-TTS (32 NFE) - Emilia (EN + ZH) - - - 2.25 2.77 0.705 0.15
MaskGCT Semantic Codec + Acoustic Codec (Wang et al., 2024) Emilia (EN + ZH) - - - 3.69 4.22 0.756 0.65
Trained on Small-Scale Dataset
VALL-E EnCodec (Défossez et al., 2023) LibriTTS 3.78 4.47 0.730 7.36 8.64 0.531 0.73
CosyVoice (Sheng et al., 2025) S3Tokenizer v1 25Hz (Du et al., 2024a) LibriTTS - - - 3.47 - - 0.45
CosyVoice 2 (Sheng et al., 2025) S3Tokenizer v2 25Hz (Du et al., 2024c) LibriTTS - - - 3.00 - - 0.45
[…]
Table 2. Subjective performance comparison on cross-sentence zero-shot speech synthesis tasks. GT duration denotes the results obtained using the ground-truth total speech duration.
[…]
Ground Truth 4.23 0
E2 TTS (32 NFE) 3.95 -0.41
F5-TTS (32 NFE) 4.04 -0.20
MaskGCT 3.98 -0.25
[…]
Table 3. Objective comparison of representative AR, NAR, and PAR models on continuation and cross-sentence zero-shot synthesis tasks. All models are trained on LibriTTS using the same speech tokenizer, detokenizer, and Transformer backbone.
ID System Language Modeling Paradigm Continuation Cross-Sentence
[…]
5. Experiments
5.1. Experimental Setup
Training Dataset We conduct experiments on the LibriTTS (Zen et al., 2019) dataset, a widely used multi-speaker English corpus containing approximately 580 hours of speech from 2,306 speakers.
Model Configurations Both stages of PALLE adopt the same model architecture, comprising: a decoder-only Transformer with 12 layers, 16 attention heads, 1024-dimensional embeddings, and 4096-dimensional feed-forward layers; a one-layer ConvNeXt V2 block with 1024-dimensional embeddings and a 2048-dimensional feed-forward layer; and a convolutional positional embedding with a kernel size of 7; totaling 177.0M parameters. For text tokenization, a 2,000-class BPE is used. Speech tokenization is carried out using the off-the-shelf S3Tokenizer [^2] https://github.com/xingchensong/S3Tokenizer from CosyVoice 2 (Du et al., 2024c) at a token rate of 25 Hz. Speech reconstruction is performed using the off-the-shelf CFM model with a built-in vocoder, also from CosyVoice 2 (Du et al., 2024c).
Training Details Both stages are trained on 8 NVIDIA TESLA V100 32GB GPUs, with a batch duration of 600 seconds per GPU. The first stage is trained for 179k steps, and the second stage is fine-tuned based on the first stage for another 87k steps. The ScaledAdam (Yao et al., 2024) optimizer and Eden (Yao et al., 2024) scheduler are used, with a peak learning rate of 0.045 for the first stage and 0.005 for the second stage.
Inference Details The first stage runs 100 steps per sample of length $T$, each generating a span of size $k=\min(\left\lfloor 0.01T\right\rfloor,N_{\text{left}})$, which corresponds to a fixed percentage of the sample length and is smaller than the span size used during training [^3] Although the model is trained with span size of $\left\lfloor 0.1T\right\rfloor$, we found the best configuration for inference is $\left\lfloor 0.01T\right\rfloor$, as detailed in Sec.  5.6.. A detailed analysis of the per-step prediction ratio is provided in Fig. 4 of the ablation study. The second stage runs 7 steps per sample using the $\gamma=0.05$ quantile of the confidence scores. Top-$p$ sampling with a small $p$ in the range of 0.2 to 0.35 yields slight performance improvements over greedy search, without introducing much randomness or additional computational cost.
5.2. Evaluation
Evaluation Settings We use the LibriSpeech (Panayotov et al., 2015) test-clean set for zero-shot TTS evaluation, ensuring no speaker overlap with training data. Following previous works (Wang et al., 2023; Chen et al., 2025), we use the same test set, which consists of segments ranging from 4 to 10 seconds, totaling 2.2 hours of data from 40 unique speakers and 1,234 samples. Following previous works (Wang et al., 2023; Chen et al., 2025), our evaluations contain two inference tasks:
  • Continuation: Given the text and the first 3 seconds of an utterance as the prompt, the model generates the remaining speech.
  • Cross-Sentence: Given a reference utterance and its transcription as the prompt, the model synthesizes speech for the target text while preserving the speaker’s characteristics.
Objective Metrics We adopt the following objective metrics to assess the naturalness, robustness, speaker similarity, and efficiency of the proposed method. For the Continuation task, we evaluate the entire utterance rather than only the continuation segment for a more comprehensive comparison.
  • WER (Word Error Rate) measures the robustness and intelligibility of synthesized speech. Neural TTS systems sometimes suffer from robustness issues, including word deletion, insertion, substitution, or generating endless silence or noise, due to attention misalignments. To evaluate these aspects, ASR models are used to transcribe the synthesized audio. WER is computed between the generated transcripts and the ground-truth text, denoted as WER-W and WER-H, with Whisper Large-v3444https://huggingface.co/openai/whisper-large-v3 (Radford et al., 2023) and the HuBERT-Large ASR model [^5] https://huggingface.co/facebook/hubert-large-ls960-ft (Hsu et al., 2021) respectively.
  • SIM-o (Speaker Similarity) measures the speaker similarity between the original speech and the synthesized speech. The state-of-the-art speaker verification model WavLM-TDNN [^6] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification#pre-trained-models (Chen et al., 2022) is employed for evaluation. The similarity score ranges from $[-1,1]$, with higher values indicating greater speaker similarity.
  • RTF (Real-Time Factor) measures the time taken to synthesize one second of speech and reflects system efficiency, especially in real-time scenarios. We report RTF on an NVIDIA TESLA A100 80G GPU, calculated from the average inference time for generating 10 seconds of speech with a batch size of 1.
Subjective Metrics We use the following subjective evaluation metrics to assess the speaker similarity and comparative quality of each synthesized speech. Twenty native English speakers with experience in speech annotation and evaluation are engaged as contributors in a crowd-sourced evaluation.
  • SMOS (Similarity Mean Opinion Score) for measuring speaker similarity between the speech prompt and the generated speech. The SMOS scale ranges from 1 to 5, in 0.5-point increments.
  • CMOS (Comparative Mean Opinion Score) for evaluating the comparative quality of the synthesized speech against the original ground-truth audio. The CMOS scale ranges from -3 (much worse than the ground truth) to 3 (much better than the ground truth), in 1-point increments.
Baseline Details We evaluate our models against several state-of-the-art zero-shot TTS systems.
  • VALL-E (Wang et al., 2023): A two-stage TTS system that utilizes both AR and NAR models to predict RVQ tokens at 75Hz from EnCodec (Défossez et al., 2023). We evaluate three versions of VALL-E: (1) the model trained on Librilight (Kahn et al., 2020), and (2) the model trained on Libriheavy (Kang et al., 2024), using performance metrics reported in (Wang et al., 2023; Chen et al., 2025) and RTF taken from (Meng et al., 2024) measured using the official checkpoint; and (3) the model trained on LibriTTS (Zen et al., 2019), which we reproduce using the open-source implementation [^7] https://github.com/lifeiteng/vall-e.
  • CosyVoice (Du et al., 2024a): A two-stage TTS system that utilizes a text-to-token AR model and a token-to-speech flow-matching model. We compare against the 25Hz version trained on LibriTTS using the official code, with results reported in (Sheng et al., 2025).
  • CosyVoice 2 (Du et al., 2024c): A two-stage TTS system that utilizes an interleaved text-speech AR model and a chunk-aware token-to-speech flow-matching model. We compare against the 25Hz version trained on LibriTTS, with results reported in (Sheng et al., 2025).
  • MaskGCT (Wang et al., 2024): A fully NAR two-stage TTS system based on mask-and-predict learning, consisting of a 695M text-to-semantic model and a 353M semantic-to-acoustic model. We use the official pre-trained checkpoint [^8] https://huggingface.co/amphion/MaskGCT, trained on 100K hours of in-the-wild Chinese and English speech data from the Emilia (He et al., 2024) dataset.
  • E2 TTS (Eskimez et al., 2024): A fully NAR TTS system based on flow matching, with a model size of  [^3] M. We use the open-source pre-trained checkpoint [^9] https://huggingface.co/SWivid/E2-TTS, trained on 100K hours of in-the-wild Chinese and English speech data from Emilia (He et al., 2024).
  • F5-TTS (Chen et al., 2024b): A fully NAR TTS system based on flow matching, with a model size of 336M. We use the official pre-trained checkpoint [^10] https://huggingface.co/SWivid/F5-TTS, trained on 100K hours of in-the-wild Chinese and English speech data from Emilia (He et al., 2024).
Figure 4. Impact of varying inference steps on objective performance metrics for cross-sentence zero-shot speech synthesis. From left to right: (a) Stage One Inference Steps on WER-H; (b) Stage One Inference Steps on SIM-o; (c) Stage Two Inference Steps on WER-H; (d) Stage Two Inference Steps on SIM-o.
[…]
Four side-by-side charts.
Figure 5. Impact of varying the total duration on objective performance for cross-sentence zero-shot speech synthesis. The left plot shows the effect of the total duration multiplier on WER-H (%), and the right plot shows the effect on SIM-o.
[…]
Two side-by-side charts. The left shows WER versus total duration multiplier. The right shows SIM versus total duration multiplier.
5.3. Evaluation Results
Objective Evaluation Table 1 presents comparisons between our proposed PALLE and the baselines in terms of robustness, similarity, and efficiency on the LibriSpeech test-clean set, where PALLE demonstrates superior performance on both continuation and cross-sentence tasks. Notably, despite being trained on small-scale data, PALLE outperforms or matches state-of-the-art systems trained on large-scale data, including F5-TTS, E2 TTS, and MaskGCT, and achieves up to 10$\times$ faster inference with an extremely low RTF of 0.06, demonstrating PAR as a powerful and efficient paradigm for codec language modeling and PALLE as a practical solution for real-time scenarios.
On the continuation task, PALLE with GT duration achieves the best WER and SIM among all systems. On the cross-sentence task, PALLE with linearly estimated duration achieves the best WER and the second-best SIM. Specifically, PALLE significantly outperforms MaskGCT and E2 TTS in intelligibility, with a 40% relative reduction in WER-W over MaskGCT, maintains competitive similarity with only a 5% relative drop, and achieves over 10$\times$ higher efficiency. Compared to the highly optimized F5-TTS, PALLE matches its robustness, surpasses it in similarity, and is 2$\times$ faster in inference speed. When compared to the LibriTTS-trained CosyVoice 2, which uses the same speech tokenizer and detokenizer, PALLE with GT duration delivers a 26% improvement in WER-W and is 7.5 times faster, both at a token rate of 25Hz.
Compared to ground-truth speech reconstructed by S3Tokenizer v2, PALLE with GT duration yields a 17% relative WER-H reduction on continuation, while PALLE with estimated duration achieves a 24% relative WER-W reduction on cross-sentence. In the latter case, it performs nearly on par with the ground-truth speech in WER-W. We also observe that PALLE with estimated duration slightly exceeds PALLE with GT duration in intelligibility while maintaining comparable similarity, suggesting the effectiveness of our BPE-level linear duration estimation in producing appropriate durations.
Subjective Evaluation The subjective evaluation uses one utterance per speaker from the LibriSpeech test-clean set, totaling 40 samples. For each sample, the preceding utterance from the official list is used as the prompt to generate the target utterance. Table 2 reports comparisons between PALLE and the baselines in terms of perceived speaker similarity and comparative quality, where PALLE with GT duration consistently outperforms all baselines across both metrics, producing speech that is virtually indistinguishable from the ground truth, with only a 0.09 gap in CMOS. PALLE with estimated duration achieves the highest CMOS, surpasses E2 TTS and MaskGCT, and performs on par with F5-TTS in SMOS.
5.4. Pseudo-Autoregressive vs. Autoregressive and Non-Autoregressive
To further assess the effectiveness of the proposed PAR modeling paradigm, we conduct a comparative study with representative AR and NAR approaches under controlled conditions. As detailed in Table 3, we train a VALL-E stage-one AR model (Sys. A) and a MaskGCT stage-one NAR model (Sys. B) on LibriTTS, using the same speech tokenizer, detokenizer, and Transformer backbone, to be compared with our PAR model (Sys. C1). Collectively, these three models serve as representative instantiations of the AR, NAR, and PAR modeling paradigms, forming a fair and isolated comparison.
PAR vs. AR Compared to the AR-based Sys. A, our PAR-based Sys. C1 achieves 36% relative reduction in WER-W and 4$\times$ speedup in inference, while maintaining comparable speaker similarity. These performance gains stem from the fundamental distinct modeling paradigms. AR models generate tokens sequentially, resulting in inference latency $\mathcal{O}(T)$, where $T=\mu D$, tied to speech token rate $\mu$ and growing linearly with duration $D$. In contrast, our PAR model generates a dynamic-length span of tokens in parallel at each step. Since each span covers a fixed proportion of the target sequence, the number of decoding steps remains constant, resulting in $\mathcal{O}(1)$ inference latency and substantial acceleration under the same token rate. Regarding accuracy, AR models rely solely on left-context for each prediction, leading to error accumulation as the sequence length increases. PAR mitigates this by predicting the entire sequence in parallel at each step and progressively committing only to the left-most span, enabling more robust and flexible temporal modeling with bidirectional context.
PAR vs. NAR On the cross-sentence task, our PAR-based Sys. C1 yields a 43% relative WER-W reduction, 2$\times$ faster inference, and marginally better speaker similarity compared to the NAR-based Sys. B. While both NAR and PAR models offer $\mathcal{O}(1)$ latency through parallel decoding, the temporally unordered generation in NAR leads to poor alignment, requiring more decoding steps $1/r^{\prime}$ with a smaller predicted proportion $r^{\prime}$ per step to maintain output quality. This results in compromised robustness, degraded prosody, and reduced efficiency in practice. PAR addresses this by introducing a soft temporal inductive bias through span-level causal ordering, preserving the natural temporal flow of speech and improving alignment and intelligibility.
5.5. Total Duration Impact Analysis
We evaluate the robustness of the stage-one and two-stage variants of PALLE under varying total duration multipliers on the cross-sentence task, where each multiplier corresponds to a different speech tempo. As shown in Fig. 5, the duration of the generated speech is scaled from 0.7$\times$ to 1.3$\times$ of the ground-truth duration. Both systems achieve the lowest WER-H at a multiplier of 1.1 and the highest SIM at 1.2, indicating that generating speech at a slightly slower pace than the ground truth improves both intelligibility and speaker similarity. One possible explanation is that, in the cross-sentence task, the speaking rate of the speech prompt may differ from that of the target. When the prompt is substantially faster, the model may struggle to maintain prosodic consistency via in-context learning, leading to degraded synthesis quality. Overall, both systems exhibit stable performance across a moderate tempo range, with near-optimal results between 0.9$\times$ and 1.3$\times$, indicating robustness to speaking rate variation.
5.6. Ablation Study
Inference Steps Fig. 4 illustrates the impact of inference steps in both stage one and stage two. In the first two plots, we vary the number of steps in stage one while keeping stage two fixed. Both curves show rapid performance gains with more steps, followed by a plateau. We select 100 steps for stage one as the optimal point to balance efficiency and performance. In the last two plots, we vary the number of steps in stage two. Although the curves exhibit some fluctuations, the overall trend improves with more steps and stabilizes after step 14. We select 7 steps for stage two as the optimal point. Although stage two involves only a few steps, it brings clear improvements across both metrics compared to stage one.
Joint Two-Stage Modeling We attempt to unify stage one and stage two into a single model via multitask finetuning, randomly performing either task per sample with equal probability. While performance slightly decreases on the continual task, WER on the cross-sentence task increases by 20%, regardless of whether ground-truth durations are used. This may be attributed to the halved parameter capacity per stage compared to the separate models.
6. Conclusions
In this paper, we propose a novel sequence generation paradigm, pseudo-autoregressive language modeling (PAR), which combines the strengths of temporal modeling from AR and parallel decoding from NAR, and generates dynamic-length spans at fixed time steps. We apply PAR to zero-shot TTS tasks, resulting in a new TTS model called PALLE. PALLE leverages a PAR model to generate initial discrete speech tokens with fixed steps for various target utterances, followed by refinement using a NAR model. Experimental results demonstrate that PALLE significantly improves intelligibility, speaker similarity, and efficiency. Looking ahead, PAR holds potential for application in broader generation domains, such as text, image, and video.

### v2

@@
 Experiments demonstrate that PALLE, trained on LibriTTS, outperforms state-of-the-art systems trained on large-scale data, including F5-TTS, E2-TTS, and MaskGCT, on the LibriSpeech test-clean set in terms of speech quality, speaker similarity, and intelligibility, while achieving up to ten times faster inference speed.
-Audio samples are available at https://anonymous-palle.github.io.
+Audio samples are available at https://microsoft.com/research/project/vall-e-x/palle.
 Zero-Shot Text-to-Speech, Pseudo-Autoregressive, Masked Language Modeling, Neural Codec Language Model
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: arXiv Preprint; April, 2025; arXiv.orgccs: Computing methodologiesccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Computing methodologies Natural language generation
-Figure 1.
-Generation procedures of three paradigms, where generation proceeds top to bottom, and each row represents one step.
-(Left) AR generates tokens step by step from left to right until the [EOS] token.
-Grouped AR can predict multiple tokens per step.
-(Middle) NAR predicts all masked positions simultaneously in a temporally unordered manner, allowing partial token retention while remasking others for regeneration.
-(Right) PAR predicts all masked positions in parallel but commits only the leftmost span at each step, enforcing generation to follow the temporal order of speech.
+journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754745isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologiesccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Computing methodologies Natural language generation
+Figure 1. Generation procedures of AR, NAR, and PAR.
 […]
@@
 1. Introduction
-Text-to-speech (TTS) synthesis has made remarkable progress, driven by generative methods (Shen et al., 2018; Li et al., 2019; Kim et al., 2021; Ren et al., 2021; Jeong et al., 2021; Wang et al., 2023) and the expansion of available data (Guo et al., 2021; Ma et al., 2024; Chen et al., 2021; Yang et al., 2024c; Kang et al., 2024; He et al., 2024).
-These innovations have enabled TTS systems to achieve human-level parity in terms of naturalness and intelligibility, in both pre-defined specific speakers (Tan et al., 2024) and zero-shot scenarios (Chen et al., 2024a).
-Modern zero-shot TTS models can be broadly divided into two paradigms: autoregressive (AR) (Wang et al., 2023; Chen et al., 2024a; Meng et al., 2024; Wang et al., 2025; Yang et al., 2024a; Du et al., 2024a, c; Peng et al., 2024; Lajszczak et al., 2024; Anastassiou et al., 2024; Kharitonov et al., 2023; Kim et al., 2024; Liu et al., 2024; Sheng et al., 2025; Guo et al., 2024) [^1] For hybrid systems like VALL-E, we regard it as AR, as primary tokens are generated autoregressively.
-and non-autoregressive (NAR) (Shen et al., 2024; Ju et al., 2024; Le et al., 2023; Eskimez et al., 2024; Liu et al., 2025; Gao et al., 2023; Chen et al., 2024b; Wang et al., 2024; Casanova et al., 2022; Jiang et al., 2023, 2024, 2025; Lee et al., 2025).
+Text-to-speech (TTS) synthesis has made remarkable progress, driven by generative methods (Shen et al., 2018; Li et al., 2019; Kim et al., 2021; Ren et al., 2021; Jeong et al., 2021; Wang et al., 2023) and the expansion of available data (Guo et al., 2021; Ma et al., 2024; Chen et al., 2021; Yang et al., 2025; Kang et al., 2024; He et al., 2024).
+These innovations have enabled TTS systems to achieve human-level parity in terms of naturalness and intelligibility, in both pre-defined specific speakers (Tan et al., 2024) and zero-shot scenarios (Chen et al., 2024).
+Modern zero-shot TTS models can be broadly divided into two paradigms: autoregressive (AR) (Wang et al., 2023; Chen et al., 2024; Meng et al., 2025; Wang et al., 2025a; Yang et al., 2024a; Sun et al., 2025; Wang et al., 2025b; Du et al., 2024a, b, 2025a; Peng et al., 2024; Lajszczak et al., 2024; Anastassiou et al., 2024; Kharitonov et al., 2023; Kim et al., 2024; Liu et al., 2024; Jia et al., 2025; Sheng et al., 2025; Guo et al., 2024; Deng et al., 2025; Zhou et al., 2025; Zhang et al., 2025; Ma et al., 2025) [^1] Hybrid systems like VALL-E are categorized by primary token generation.
+and non-autoregressive (NAR) (Shen et al., 2024; Ju et al., 2024; Le et al., 2023; Eskimez et al., 2024; Liu et al., 2025; Gao et al., 2023; Chen et al., 2025a; Wang et al., 2024; Casanova et al., 2022; Jiang et al., 2023, 2024, 2025; Lee et al., 2025; Zhu et al., 2025).
 AR models sequentially generate speech from left to right, conditioning each step on previous outputs.
@@
 Specifically, the fixed prediction size per step results in slow inference, with generation steps tightly coupled to the frame rate of output units.
-Additionally, the strictly unidirectional modeling restricts access to left context only and lacks future or global information, leading to exposure bias and posing challenges to robustness (Song et al., 2024a; Han et al., 2024; Du et al., 2024b).
-Techniques like grouped modeling (Chen et al., 2024a) and speculative decoding (Li et al., 2024) improve efficiency, but generation steps still scale linearly with synthesized speech duration.
+Additionally, the strictly unidirectional modeling restricts access to left context only and lacks future or global information, leading to exposure bias and posing challenges to robustness (Song et al., 2025; Han et al., 2024; Du et al., 2025b).
+Techniques like grouped modeling (Chen et al., 2024) and speculative decoding (Li et al., 2025) improve efficiency, but generation steps still scale linearly with synthesized speech duration.
 In contrast, NAR models enable fast inference through parallel generation.
-NAR TTS methods leverage denoising diffusion (Ju et al., 2024), flow matching (Eskimez et al., 2024; Chen et al., 2024b), GAN (Jiang et al., 2024), and masked generative modeling (Wang et al., 2024), drawing inspiration from advances in image (Chang et al., 2022) and language (Devlin et al., 2019) generation.
+NAR TTS methods leverage denoising diffusion (Ju et al., 2024), flow matching (Eskimez et al., 2024; Chen et al., 2025a), GAN (Jiang et al., 2024), and masked generative modeling (Wang et al., 2024), drawing inspiration from advances in image (Chang et al., 2022) and language (Devlin et al., 2019) generation.
 Recent advances (Eskimez et al., 2024; Wang et al., 2024) demonstrate that text-to-speech alignment can be achieved without phoneme-level duration prediction or explicit alignment supervision.
-E2-TTS (Eskimez et al., 2024) and F5-TTS (Chen et al., 2024b) simply pad the character sequence with filler tokens to match the linearly predicted or user-specified speech length.
+E2-TTS (Eskimez et al., 2024) and F5-TTS (Chen et al., 2025a) simply pad the character sequence with filler tokens to match the linearly predicted or user-specified speech length.
 MaskGCT (Wang et al., 2024) concatenates the tokenized text and speech sequences, adopts mask-and-predict training, and performs iterative parallel decoding during inference.
-However, despite their efficiency and simplicity, these methods lack temporal modeling and generate speech in a temporally unordered manner, resulting in slow convergence (Eskimez et al., 2024; Chen et al., 2024b), alignment inaccuracies (Eskimez et al., 2024), and degraded intelligibility (Wang et al., 2024).
+However, despite their efficiency and simplicity, these methods lack temporal modeling and generate speech in a temporally unordered manner, resulting in slow convergence (Eskimez et al., 2024; Chen et al., 2025a), alignment inaccuracies (Eskimez et al., 2024), and degraded intelligibility (Wang et al., 2024).
 To combine the strengths of AR and NAR paradigms while mitigating their limitations, we propose a novel pseudo-autoregressive (PAR) codec language modeling paradigm that unifies AR and NAR modeling by combining temporal modeling from AR with parallel generation from NAR.
@@
 Local spans are generated simultaneously, with earlier ones committed before later ones, creating a span-level temporal flow that aligns with the structure of natural speech.
-Unlike grouped AR methods (Chen et al., 2024a), which generate fixed-length spans with dynamic time steps, PAR generates dynamic-length spans with fixed time steps.
+Unlike grouped AR methods (Chen et al., 2024), which generate fixed-length spans with dynamic time steps, PAR generates dynamic-length spans with fixed time steps.
 Building on PAR, we propose PALLE, a two-stage TTS system that performs PAR generation followed by confidence-guided refinement.
@@
   • We achieve a breakthrough in zero-shot TTS performance and efficiency, with the small-scale trained PALLE outperforming large-scale state-of-the-art systems while enabling up to 10$\times$ faster inference.
-2. Related work
-2.1. Zero-Shot TTS
-Zero-shot TTS enables speech synthesis for unseen speakers by capturing the timbre, prosody, and style from short speech prompts.
-Benefiting from in-context learning (ICL), the performance of zero-shot TTS has improved significantly, reaching human-parity quality (Ju et al., 2024; Chen et al., 2024a).
-Recent approaches primarily fall into three categories: codec language modeling (Wang et al., 2023; Chen et al., 2024a; Wang et al., 2024), diffusion/flow-based methods (Shen et al., 2024; Ju et al., 2024; Le et al., 2023; Eskimez et al., 2024; Chen et al., 2024b), and hybrid approaches (Du et al., 2024a, c).
-Hybrid systems involve a text-to-codec language model with a codec-to-mel model based on diffusion or flow matching, disentangling complex speech components to improve zero-shot generalization.
 Figure 2.
@@
 An overview diagram of PALLE.
+2. Related work
+2.1. Zero-Shot TTS
+Zero-shot TTS enables speech synthesis for unseen speakers by capturing the timbre, prosody, and style from short speech prompts.
+Benefiting from in-context learning (ICL), the performance of zero-shot TTS has improved significantly, reaching human-parity quality (Ju et al., 2024; Chen et al., 2024).
+Recent approaches primarily fall into three categories: codec language modeling (Wang et al., 2023; Chen et al., 2024; Wang et al., 2024), diffusion/flow-based methods (Shen et al., 2024; Ju et al., 2024; Le et al., 2023; Eskimez et al., 2024; Chen et al., 2025a), and hybrid approaches (Du et al., 2024a, b).
+Hybrid systems involve a text-to-codec language model with a codec-to-mel model based on diffusion or flow matching, disentangling complex speech components to improve zero-shot generalization.
 2.2. Discrete Speech Representation
 Discrete speech representations (Yang et al., 2024b) are typically categorized into acoustic tokens and semantic tokens.
-Acoustic tokens, produced by neural codec models (Défossez et al., 2023; Kumar et al., 2023; Zeghidour et al., 2022; Ji et al., 2025; Parker et al., 2024; Wu et al., 2024; Du et al., 2024d), aim to reconstruct speech with high fidelity, capturing fine-grained acoustic details.
+Acoustic tokens, produced by neural codec models (Défossez et al., 2023; Kumar et al., 2023; Zeghidour et al., 2022; Ji et al., 2025; Parker et al., 2024; Wu et al., 2024; Du et al., 2024c), aim to reconstruct speech with high fidelity, capturing fine-grained acoustic details.
 In contrast, semantic tokens are quantized from embeddings of self-supervised learning (SSL) models (Baevski et al., 2020a, b; Hsu et al., 2021; Chen et al., 2022), which are trained for discrimination or masked prediction, and preserve the linguistic content of speech.
-Recent methods have explored supervised semantic tokens (Du et al., 2024a, c), trained with automatic speech recognition (ASR) loss (Graves et al., 2006, 2013), making them more robust to noise and better suited for language modeling.
+Recent methods have explored supervised semantic tokens (Du et al., 2024a, b), trained with automatic speech recognition (ASR) loss (Graves et al., 2006, 2013), making them more robust to noise and better suited for language modeling.
 2.3. Neural Codec Language Modeling
@@
 Two main paradigms have been explored: AR and NAR modeling.
-AR models (Wang et al., 2023; Chen et al., 2024a; Yang et al., 2024a; Du et al., 2024a, c; Kharitonov et al., 2023; Lajszczak et al., 2024; Sheng et al., 2025) generate tokens sequentially from left to right, providing strong temporal grounding but suffering from slow inference and limited context modeling.
+AR models (Wang et al., 2023; Chen et al., 2024; Yang et al., 2024a; Du et al., 2024a, b; Kharitonov et al., 2023; Lajszczak et al., 2024; Sheng et al., 2025) generate tokens sequentially from left to right, providing strong temporal grounding but suffering from slow inference and limited context modeling.
 SyncSpeech (Sheng et al., 2025) is an AR masked generative model that incorporates temporal modeling via a causal attention mask and token-level duration prediction, aiming for streaming generation from partial text input.
@@
 MaskGCT (Wang et al., 2024) predicts the total speech length and performs temporally unordered prediction, relying solely on confidence-based scheduling, which may cause robustness issues in high-entropy regions such as fast transitions or expressive prosody.
+Figure 3.
+Illustration of the two-stage PALLE framework.
+(a) The shared architecture: a bidirectional masked generative transformer.
+(b) Stage one: the PAR model predicts all token positions in parallel but retains only the leftmost span at each step.
+(c) Stage two: the NAR model refines the initial speech tokens, where low-confidence tokens are re-masked and re-generated using contextual information.
+[…]
+An illustration diagram of PALLE framework.
 3. Pseudo-Autoregressive Language Modeling
@@
 During training, the model is optimized to maximize the likelihood of the speech tokens $\mathbf{c}$ conditioned on the text content $\mathbf{x}$ using cross-entropy loss.
-At each training step, a contiguous span mask is applied to $\mathbf{c}$ from a random starting position to the end:
+At each step, a contiguous span mask is applied to $\mathbf{c}$ from a random starting position to the end:
 (1) $\mathbf{m}=[\underbrace{0,0,\ldots,0}_{T-\ell},\underbrace{1,1,\ldots,1}_{\ell}],$
@@
 Note that all self-attention layers use full attention.
-During inference, given a tokenized text $\mathbf{x}^{\text{gen}}$ to be synthesized, a speech prompt $\mathbf{y}^{\text{ref}}$ from an unseen speaker, and its corresponding tokenized text $\mathbf{x}^{\text{ref}}$, the speech tokens $\mathbf{c}^{\text{ref}}=[c^{\prime}_{0},c^{\prime}_{1},\ldots,c^{\prime}_{T^{\text{ref}}-1}]$ of length $T^{\text{ref}}$.
+During inference, given a tokenized text $\mathbf{x}^{\text{gen}}$ to be synthesized, a speech prompt $\mathbf{y}^{\text{ref}}$ from an unseen speaker, and its corresponding tokenized text $\mathbf{x}^{\text{ref}}$, the speech tokens $\mathbf{c}^{\text{ref}}=[c^{\prime}_{0},c^{\prime}_{1},\ldots,c^{\prime}_{T^{\text{ref}}-1}]$ of length $T^{\text{ref}}$ are extracted.
 The target speech token length $T^{\text{gen}}$ can be determined arbitrarily or estimated.
@@
 where the appended zeros serve as placeholders for the speech tokens to be generated.
-At each decoding step $t$, the model predicts all tokens in parallel, but retains only the leftmost $k^{\prime}=\min(\lfloor r^{\prime}T^{\text{gen}}\rfloor,N_{\text{left}})$ tokens, where $r^{\prime}\in(0,1)$ is a fixed ratio and $N_{\text{left}}$ is the number of tokens yet to be generated.
+At each step $t$, the model predicts all tokens in parallel, but retains only the leftmost $k^{\prime}=\min(\lfloor r^{\prime}T^{\text{gen}}\rfloor,N_{\text{left}})$ tokens, where $r^{\prime}\in(0,1)$ is a fixed ratio and $N_{\text{left}}$ is the number of tokens yet to be generated.
 Predictions are conditioned on the text prompt $\mathbf{x}^{\text{ref}}$, the target text $\mathbf{x}^{\text{gen}}$, and the current extended speech token sequence $\mathbf{c}^{\text{ext}}_{(t)}$.
@@
 It accommodates both temporal concatenation, as in MaskGCT (Wang et al., 2024), and feature-dimension fusion, where text and speech tokens are first padded to the same length, then independently embedded and concatenated along the feature dimension, as in E2 TTS (Eskimez et al., 2024).
-Figure 3.
-Illustration of the two-stage PALLE framework.
-(a) The shared architecture: a bidirectional masked generative transformer.
-(b) Stage one: the model predicts all token positions in parallel but retains only the leftmost span at each step.
-(c) Stage two: the model refines the initial speech tokens, where low-confidence tokens are re-masked and regenerated using contextual information.
-[…]
-An illustration diagram of PALLE framework.
 4. PALLE
@@
 This model is a fine-tuned version of the SenseVoice-Large (An et al., 2024) ASR model, trained on a large multilingual speech dataset for robust speech understanding.
-By leveraging ASR loss during training, the S3Tokenizer extracts semantic information, enabling implicit denoising and speaker disentanglement (Song et al., 2024b).
-We adopt the pretrained conditional flow matching (CFM) model from CosyVoice 2 (Du et al., 2024c) to reconstruct mel spectrograms from discrete speech tokens.
-The generated mel spectrogram is then converted into the waveform using the pretrained HiFi-GAN (Kong et al., 2020) vocoder from CosyVoice 2 (Du et al., 2024c).
+By leveraging ASR loss during training, the S3Tokenizer extracts semantic information, enabling implicit denoising and speaker disentanglement (Song et al., 2024).
+We adopt the pretrained conditional flow matching (CFM) model from CosyVoice 2 (Du et al., 2024b) to reconstruct mel spectrograms from discrete speech tokens.
+The generated mel spectrogram is then converted into the waveform using the pretrained HiFi-GAN (Kong et al., 2020) vocoder from CosyVoice 2 (Du et al., 2024b).
 4.2. Training: Conditional Codec Masked Language Modeling
@@
 We reuse the symbol $\mathbf{x}^{\text{gen}}=[x^{g}_{0},x^{g}_{1},\ldots,x^{g}_{L^{\text{gen}}-1}]$ of length $L^{\text{gen}}$ to represent the tokenized text to be synthesized, $\mathbf{y}^{\text{ref}}$ as the speech prompt, $\mathbf{x}^{\text{ref}}=[x^{r}_{0},x^{r}_{1},\ldots,x^{r}_{L^{\text{ref}}-1}]$ of length $L^{\text{ref}}$ as the tokenized text prompt, and $\mathbf{c}^{\text{ref}}=[c^{r}_{0},c^{r}_{1},\ldots,c^{r}_{T^{\text{ref}}-1}]$ of length $T^{\text{ref}}$ as the corresponding speech tokens from $\mathbf{y}^{\text{ref}}$.
-Stage One The target speech token length $T^{\text{gen}}$ is either predefined or estimated linearly as in (Chen et al., 2024b):
+Stage One The target speech token length $T^{\text{gen}}$ is either predefined or estimated linearly as in (Chen et al., 2025a):
 (11) $T^{\text{gen}}=T^{\text{ref}}\times(1+\frac{L^{\text{gen}}}{L^{\text{ref}}}).$
@@
 Table 1.
-Objective performance comparison on continuation and cross-sentence zero-shot speech synthesis tasks.
+Objective performance comparison on continuation and cross-sentence tasks.
 Bold highlights the best result, while underlined marks the second-best.
 Metrics not reported in the original papers are calculated using the open-source checkpoints.
-GT duration denotes the results using the ground-truth total speech duration.
+GT duration uses ground-truth total speech duration.
 For Continuation task, NAR systems without GT duration are excluded, as linear duration estimation requires forced alignment to obtain the transcription of the first three seconds of speech.
@@
 VALL-E (Wang et al., 2023) EnCodec (Défossez et al., 2023) Librilight - 3.80 0.773 - 5.90 0.633 0.73
-VALL-E  (Chen et al., 2025) EnCodec (Défossez et al., 2023) Libriheavy - 3.00 0.770 - 3.80 0.630 0.73
+VALL-E  (Chen et al., 2025b) EnCodec (Défossez et al., 2023) Libriheavy - 3.00 0.770 - 3.80 0.630 0.73
 E2 TTS (32 NFE) - Emilia (EN + ZH) - - - 2.55 2.92 0.756 0.68
@@
 CosyVoice (Sheng et al., 2025) S3Tokenizer v1 25Hz (Du et al., 2024a) LibriTTS - - - 3.47 - - 0.45
-CosyVoice 2 (Sheng et al., 2025) S3Tokenizer v2 25Hz (Du et al., 2024c) LibriTTS - - - 3.00 - - 0.45
+CosyVoice 2 (Sheng et al., 2025) S3Tokenizer v2 25Hz (Du et al., 2024b) LibriTTS - - - 3.00 - - 0.45
 […]
 […]
-Table 2. Subjective performance comparison on cross-sentence zero-shot speech synthesis tasks. GT duration denotes the results obtained using the ground-truth total speech duration.
+Table 2.
+Subjective performance comparison on cross-sentence zero-shot speech synthesis tasks.
+GT duration denotes the results obtained using the ground-truth total speech duration.
+CI (95%) shows 95% confidence intervals.
+p-value is obtained from the Wilcoxon signed-rank test to evaluate statistical significance.
 […]
-Ground Truth 4.23 0
-E2 TTS (32 NFE) 3.95 -0.41
-F5-TTS (32 NFE) 4.04 -0.20
-MaskGCT 3.98 -0.25
+Ground Truth 4.23 0 0.18 -
+E2 TTS (32 NFE) 3.95 -0.41 0.20 0.031
+F5-TTS (32 NFE) 4.04 -0.20 0.18 0.033
+MaskGCT 3.98 -0.25 0.17 0.025
 […]
@@
 For text tokenization, a 2,000-class BPE is used.
-Speech tokenization is carried out using the off-the-shelf S3Tokenizer [^2] https://github.com/xingchensong/S3Tokenizer from CosyVoice 2 (Du et al., 2024c) at a token rate of 25 Hz.
-Speech reconstruction is performed using the off-the-shelf CFM model with a built-in vocoder, also from CosyVoice 2 (Du et al., 2024c).
+Speech tokenization is carried out using the off-the-shelf S3Tokenizer [^2] https://github.com/xingchensong/S3Tokenizer from CosyVoice 2 (Du et al., 2024b) at a token rate of 25 Hz.
+Speech reconstruction is performed using the off-the-shelf CFM model with a built-in vocoder, also from CosyVoice 2 (Du et al., 2024b).
 Training Details Both stages are trained on 8 NVIDIA TESLA V100 32GB GPUs, with a batch duration of 600 seconds per GPU.
@@
 Inference Details The first stage runs 100 steps per sample of length $T$, each generating a span of size $k=\min(\left\lfloor 0.01T\right\rfloor,N_{\text{left}})$, which corresponds to a fixed percentage of the sample length and is smaller than the span size used during training [^3] Although the model is trained with span size of $\left\lfloor 0.1T\right\rfloor$, we found the best configuration for inference is $\left\lfloor 0.01T\right\rfloor$, as detailed in Sec.
-5.6..
+5.5..
 A detailed analysis of the per-step prediction ratio is provided in Fig.
@@
 Evaluation Settings We use the LibriSpeech (Panayotov et al., 2015) test-clean set for zero-shot TTS evaluation, ensuring no speaker overlap with training data.
-Following previous works (Wang et al., 2023; Chen et al., 2025), we use the same test set, which consists of segments ranging from 4 to 10 seconds, totaling 2.2 hours of data from 40 unique speakers and 1,234 samples.
-Following previous works (Wang et al., 2023; Chen et al., 2025), our evaluations contain two inference tasks:
+Following previous works (Wang et al., 2023; Chen et al., 2025b), we use the same test set, which consists of segments ranging from 4 to 10 seconds, totaling 2.2 hours of data from 40 unique speakers and 1,234 samples.
+Our evaluations contain two inference tasks:
   • Continuation: Given the text and the first 3 seconds of an utterance as the prompt, the model generates the remaining speech.
-  • Cross-Sentence: Given a reference utterance and its transcription as the prompt, the model synthesizes speech for the target text while preserving the speaker’s characteristics.
+  • Cross-Sentence: Given a reference utterance and its transcription as the prompt, the model synthesizes speech for the target text.
 Objective Metrics We adopt the following objective metrics to assess the naturalness, robustness, speaker similarity, and efficiency of the proposed method.
@@
   • WER (Word Error Rate) measures the robustness and intelligibility of synthesized speech.
-Neural TTS systems sometimes suffer from robustness issues, including word deletion, insertion, substitution, or generating endless silence or noise, due to attention misalignments.
-To evaluate these aspects, ASR models are used to transcribe the synthesized audio.
-WER is computed between the generated transcripts and the ground-truth text, denoted as WER-W and WER-H, with Whisper Large-v3444https://huggingface.co/openai/whisper-large-v3 (Radford et al., 2023) and the HuBERT-Large ASR model [^5] https://huggingface.co/facebook/hubert-large-ls960-ft (Hsu et al., 2021) respectively.
+WER is computed between ASR transcription of synthesized audio and the ground-truth text, denoted as WER-W and WER-H, with Whisper Large-v3444https://huggingface.co/openai/whisper-large-v3 (Radford et al., 2023) and the HuBERT-Large ASR model [^5] https://huggingface.co/facebook/hubert-large-ls960-ft (Hsu et al., 2021) respectively.
   • SIM-o (Speaker Similarity) measures the speaker similarity between the original speech and the synthesized speech.
@@
 Baseline Details We evaluate our models against several state-of-the-art zero-shot TTS systems.
-  • VALL-E (Wang et al., 2023): A two-stage TTS system that utilizes both AR and NAR models to predict RVQ tokens at 75Hz from EnCodec (Défossez et al., 2023).
-We evaluate three versions of VALL-E: (1) the model trained on Librilight (Kahn et al., 2020), and (2) the model trained on Libriheavy (Kang et al., 2024), using performance metrics reported in (Wang et al., 2023; Chen et al., 2025) and RTF taken from (Meng et al., 2024) measured using the official checkpoint; and (3) the model trained on LibriTTS (Zen et al., 2019), which we reproduce using the open-source implementation [^7] https://github.com/lifeiteng/vall-e.
-  • CosyVoice (Du et al., 2024a): A two-stage TTS system that utilizes a text-to-token AR model and a token-to-speech flow-matching model.
-We compare against the 25Hz version trained on LibriTTS using the official code, with results reported in (Sheng et al., 2025).
-  • CosyVoice 2 (Du et al., 2024c): A two-stage TTS system that utilizes an interleaved text-speech AR model and a chunk-aware token-to-speech flow-matching model.
-We compare against the 25Hz version trained on LibriTTS, with results reported in (Sheng et al., 2025).
-  • MaskGCT (Wang et al., 2024): A fully NAR two-stage TTS system based on mask-and-predict learning, consisting of a 695M text-to-semantic model and a 353M semantic-to-acoustic model.
-We use the official pre-trained checkpoint [^8] https://huggingface.co/amphion/MaskGCT, trained on 100K hours of in-the-wild Chinese and English speech data from the Emilia (He et al., 2024) dataset.
-  • E2 TTS (Eskimez et al., 2024): A fully NAR TTS system based on flow matching, with a model size of  [^3] M.
-We use the open-source pre-trained checkpoint [^9] https://huggingface.co/SWivid/E2-TTS, trained on 100K hours of in-the-wild Chinese and English speech data from Emilia (He et al., 2024).
-  • F5-TTS (Chen et al., 2024b): A fully NAR TTS system based on flow matching, with a model size of 336M.
-We use the official pre-trained checkpoint [^10] https://huggingface.co/SWivid/F5-TTS, trained on 100K hours of in-the-wild Chinese and English speech data from Emilia (He et al., 2024).
+  • VALL-E (Wang et al., 2023): A two-stage AR+NAR system predicting 75Hz EnCodec RVQ tokens (Défossez et al., 2023).
+We evaluate three variants trained on Librilight (Kahn et al., 2020), Libriheavy (Kang et al., 2024) and LibriTTS (Zen et al., 2019), with RTF from (Meng et al., 2025) using the official checkpoint.
+  • CosyVoice (Du et al., 2024a): A two-stage AR+flow-matching system. We evaluate the 25Hz LibriTTS version, with results from (Sheng et al., 2025).
+  • CosyVoice 2 (Du et al., 2024b): A two-stage interleaved AR+chunk-aware flow-matching system. We evaluate the 25Hz LibriTTS version, with results from (Sheng et al., 2025).
+  • MaskGCT (Wang et al., 2024): A two-stage NAR mask-and-predict system with 695M text-to-semantic and 353M semantic-to-acoustic models.
+We evaluate the official checkpoint [^7] https://huggingface.co/amphion/MaskGCT trained on 100K hours of Chinese and English speech from Emilia (He et al., 2024).
+  • E2 TTS (Eskimez et al., 2024) & F5-TTS (Chen et al., 2025a): Fully NAR flow-matching systems with  [^3] M and 336M parameters.
+We evaluate open-source checkpoints [^8] https://huggingface.co/SWivid/E2-TTS [^9] https://huggingface.co/SWivid/F5-TTS trained on 100K hours of bilingual Emilia (He et al., 2024).
 Figure 4.
@@
 PAR addresses this by introducing a soft temporal inductive bias through span-level causal ordering, preserving the natural temporal flow of speech and improving alignment and intelligibility.
-5.5. Total Duration Impact Analysis
-We evaluate the robustness of the stage-one and two-stage variants of PALLE under varying total duration multipliers on the cross-sentence task, where each multiplier corresponds to a different speech tempo.
+5.5. Ablation Study
+Total Duration We evaluate the robustness of the stage-one and two-stage variants of PALLE under varying total duration multipliers on the cross-sentence task, where each multiplier corresponds to a different speech tempo.
 As shown in Fig.
@@
 Overall, both systems exhibit stable performance across a moderate tempo range, with near-optimal results between 0.9$\times$ and 1.3$\times$, indicating robustness to speaking rate variation.
-5.6. Ablation Study
 Inference Steps Fig.
@@
 Although stage two involves only a few steps, it brings clear improvements across both metrics compared to stage one.
-Joint Two-Stage Modeling We attempt to unify stage one and stage two into a single model via multitask finetuning, randomly performing either task per sample with equal probability.
-While performance slightly decreases on the continual task, WER on the cross-sentence task increases by 20%, regardless of whether ground-truth durations are used.
+Joint Two-Stage Modeling We attempt to unify stage one and stage two into a single model via multitask finetuning, randomly performing one task per sample.
+However, stage two loss degrades stage one performance.
+While performance slightly decreases on the continuation task, WER on the cross-sentence task increases by 20%, regardless of whether ground-truth durations are used.
 This may be attributed to the halved parameter capacity per stage compared to the separate models.
@@
 Experimental results demonstrate that PALLE significantly improves intelligibility, speaker similarity, and efficiency.
-Looking ahead, PAR holds potential for application in broader generation domains, such as text, image, and video.
+Looking ahead, PAR holds potential for broader applicability in generation domains, such as text, image, and video.
+7. Acknowledgments
+This work was supported by the National Natural Science Foundation of China (No.
+62206171 and No.
+U23B2018), Shanghai Municipal Science and Technology Major Project under Grant 2021SHZDZX0102 and Yangtze River Delta Science and Technology Innovation Community Joint Research Project (2024CSJGG01100).